OpenStudy (anonymous):

can you explain 26n+15=32n+3 . im stuck on this one

OpenStudy (anonymous):

Well we want the n on one side and the constants on the other. 15 -3 = 32n - 26n 12 = 6n now we divide 12/6 = 2 = n